# Existence of All Wilton Ripples of the Kawahara Equation

### v1

@@ 2.3.1 Case 1: $K=2$
$|\mathcal{R}_{\omega,2}|$ $\leq 2\big{(}|a|+|b|\big{)}||R_{3}(x;a,b,c_{r})||_{\textrm{L}^{1}(\mathbb{T})},$
$\leq 2\sqrt{2\pi}(|a|+|b|)||R_{3}(x;a,b,c_{r})||_{\textrm{L}^{2}(\mathbb{T})},$
$\leq 2\sqrt{2\pi}(|a|+|b|)||R_{3}(x;a,b,c_{r})||_{\textrm{H}^{4}(\mathbb{T})},$
@@
~$\tilde{\mathcal{B}}(a,b,c_{r})$ $=\begin{pmatrix}ab+ac_{r}-\frac{5}{4}a^{3}-\frac{11}{8}ab^{2}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}\\ \frac{1}{2}a^{2}+bc_{r}-\frac{11}{8}a^{2}b-\frac{91}{72}b^{3}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}\end{pmatrix}=0.$ (52)
@@
$b=\frac{a}{\sqrt{2}}\tilde{b},\quad c_{r}=-\frac{a}{\sqrt{2}}\tilde{c_{r}},\quad\textrm{and}\quad\tilde{\tilde{\mathcal{B}}}(a,\tilde{b},\tilde{c_{r}})=\frac{1}{a^{2}}\tilde{\mathcal{B}}\left(a,\frac{a}{\sqrt{2}}\tilde{b},-\frac{a}{\sqrt{2}}\tilde{c_{r}}\right).$ (53)
@@
From (54), we have that ~~$\tilde{\tilde{\mathcal{B}}}(0,\pm 1,\pm 1)=0$. The corresponding Jacobian evaluated at $(0,\pm 1,\pm 1)$ is
~~$\frac{\partial\tilde{\tilde{\mathcal{B}}}}{\partial(\tilde{b},\tilde{c_{r}})}(0,\pm 1,\pm 1)=\begin{pmatrix}\frac{1}{\sqrt{2}}&-\frac{1}{\sqrt{2}}\\ \mp 1&\mp 1\end{pmatrix},$ (55)
which is invertible in both cases. The implicit function theorem guarantees the existence and uniqueness of
@@
such that, for sufficiently small $a$, we have (i) ~~$\tilde{\tilde{\mathcal{B}}}(a,\tilde{b}^{\pm}(a),\tilde{c_{r}}^{\pm}(a))=0$, (ii) $\tilde{b}^{\pm}(0)=\pm 1$ and $\tilde{c_{r}}^{\pm}(0)=\pm 1$, and (iii) $\tilde{b}^{\pm}(a)$ and $\tilde{c_{r}}^{\pm}(a)$ are real analytic in $a$.
@@
$c=c_{0}-\frac{a}{\sqrt{2}}\tilde{c_{r}}^{\pm}(a)=\frac{4}{5}-\frac{a}{\sqrt{2}}\tilde{c_{r}}^{\pm}(a),$ (57)
@@ 2.3.2 Case 2: $K=3$
~$\tilde{\mathcal{B}}(a,b,c_{r})$ $=\begin{pmatrix}ac_{r}-\frac{7}{9}a^{3}+a^{2}b-\frac{34}{63}ab^{2}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}\\ bc_{r}+\frac{1}{3}a^{3}-\frac{34}{63}a^{2}b-\frac{211}{189}b^{3}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}\end{pmatrix}=0.$ (62)
@@
$b=a\tilde{b},\quad c_{r}=a^{2}\tilde{c_{r}},\quad\textrm{and}\quad\tilde{\tilde{\mathcal{B}}}(a,\tilde{b},\tilde{c_{r}})=\frac{1}{a^{3}}\tilde{B}(a,a\tilde{b},a^{2}\tilde{c_{r}}),$ (63)
@@
~~$\tilde{\tilde{\mathcal{B}}}(a,\tilde{b},\tilde{c_{r}})=\begin{pmatrix}-\frac{7}{9}+\tilde{c_{r}}+\tilde{b}-\frac{34}{63}\tilde{b}^{2}+\mathcal{O}\left(|a|\left(1+|\tilde{b}|+|a|^{2}|\tilde{c_{r}}|\right)^{4}\right)\\ \frac{1}{3}-\frac{34}{63}\tilde{b}+\tilde{b}\tilde{c_{r}}-\frac{211}{189}\tilde{b}^{3}+\mathcal{O}\left(|a|\left(1+|\tilde{b}|+|a|^{2}|\tilde{c_{r}}|\right)^{4}\right)\end{pmatrix}=0.$ (64)
@@
~~$\tilde{\tilde{\mathcal{B}}}\left(0,\tilde{b}_{0}^{(\sigma)},\tilde{c_{r,0}}^{(\sigma)}\right)=0,$ (68)
@@
~~$\frac{\partial\tilde{\tilde{\mathcal{B}}}}{\partial(\tilde{b},\tilde{c_{r}})}(0,\tilde{b}_{0}^{(\sigma)},\tilde{c_{r,0}}^{(\sigma)})=\begin{pmatrix}1&1\\ -\frac{34}{63}&\tilde{b}_{0}^{(\sigma)}\end{pmatrix},$ (69)
which are invertible since $\tilde{b}_{0}^{(\sigma)}\neq-34/63$ for all $\sigma$ by contradiction. The implicit function theorem then gives the existence and uniqueness of
@@
for each $\sigma\in\{1,2,3\}$ and sufficiently small $a$ such that (i) ~~$\tilde{\tilde{\mathcal{B}}}(a,\tilde{b}^{(\sigma)}(a),\tilde{c_{r}}^{(\sigma)}(a))=0$, (ii) $\tilde{b}^{(\sigma)}(0)=\tilde{b}_{0}^{(\sigma)}$ and $\tilde{c_{r}}^{(\sigma)}(0)=\tilde{c_{r,0}}^{(\sigma)}$, and (iii) $\tilde{b}^{(\sigma)}(a)$ and $\tilde{c_{r}}^{(\sigma)}(a)$ are real analytic in $a$.
@@
$c=c_{0}+a^{2}\tilde{c_{r}}^{(\sigma)}(a)=\frac{9}{10}+a^{2}\tilde{c_{r}}^{(\sigma)}(a),$ (71)
@@ 2.3.3 Case 3: $K\geq 4$
~$\tilde{\mathcal{B}}(a,b,c_{r})$ $=\begin{pmatrix}ac_{r}+v_{3,0,0}(K)a^{3}+v_{1,2,0}(K)ab^{2}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}\\ bc_{r}+w_{2,1,0}(K)a^{2}b+w_{0,3,0}(K)b^{3}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}\end{pmatrix}=0.$ (77)
@@
~~$\tilde{\tilde{\mathcal{B}}}(a,\tilde{b},\tilde{c_{r}})=\begin{pmatrix}v_{3,0,0}(K)+\tilde{c_{r}}+v_{1,2,0}(K)\tilde{b}^{2}+\mathcal{O}\left(|a|\left(1+|\tilde{b}|+|a|^{2}|\tilde{c_{r}}|\right)^{4}\right)\\ w_{2,1,0}(K)\tilde{b}+\tilde{c_{r}}\tilde{b}+w_{0,3,0}(K)\tilde{b}^{3}+\mathcal{O}\left(|a|\left(1+|\tilde{b}|+|a|^{2}|\tilde{c_{r}}|\right)^{4}\right)\end{pmatrix}=0.$ (78)
@@
~~$\tilde{\tilde{\mathcal{B}}}(0,0,-v_{3,0,0}(K))=0,$ (81)
@@
~~$\frac{\partial\tilde{\tilde{\mathcal{B}}}}{\partial(\tilde{b},\tilde{c_{r}})}(0,0,-v_{3,0,0}(K))=\begin{pmatrix}0&1\\ w_{2,1,0}(K)-v_{3,0,0}(K)&0\end{pmatrix},$ (82)
@@
Thus, (82) is invertible, and the implicit function theorem gives us real analytic functions $\tilde{b}(a)$ and $\tilde{c_{r}}(a)$ such that ~~$\tilde{\tilde{\mathcal{B}}}(a,\tilde{b}(a),\tilde{c_{r}}(a))=0$, $\tilde{b}(0)=0$, and $\tilde{c_{r}}(0)=-v_{3,0,0}(K)$.
@@
$c=c_{0}+a^{2}\tilde{c_{r}}(a)=\frac{K^{2}}{K^{2}+1}+a^{2}\tilde{c_{r}}(a),$ (84)
@@
As before, one can show that $u(x;a)$ is real analytic in $a$ with values in $\textrm{H}^{\infty}_{\textrm{per},\textrm{even}}(\mathbb{T})$.
Unlike in previous cases, we cannot yet claim that (85) are Wilton ripple solutions. In particular, we must ensure that $\tilde{b}(a)$ does not identically vanish for all sufficiently small $a$. This is not so easy to verify. However, given we have rigorously proven that $u(x;a)$ and $c(a)$ are real analytic in $a$, one can return to (5), substitute in the appropriate expansions for $u(x;a)$ and $c(a)$, and apply a Poincaré-Linstedt perturbation method to obtain increasingly higher-order coefficients of these expansions. These calculations have been carried out in great detail by [14], where analyticity was assumed a priori. In Lemma 5 of this work, it was shown that $\mathcal{F}_{K}(u(x;a))\neq 0$ and, in fact, $\mathcal{F}_{K}(u(x;a))=\mathcal{O}(a^{K-2})$ as $a\rightarrow 0$. Hence, we conclude $\tilde{b}(a)\not\equiv 0$ and, more specifically, $\tilde{b}(a)=\mathcal{O}(a^{K-3})$ as $a\rightarrow 0$, completing the proof of the final case of Theorem 1.
3 Conclusions
Wilton ripples represent a special class of small-amplitude, periodic, traveling-wave solutions that bifurcate from two co-propagating cosine waves. In general, these solutions represent important resonances in the study of water waves [24] as well as in magnetohydrodynamics [26]. Current existence proofs of Wilton ripples [2]-[3], [12, 16], and [19]-[20] mostly consider the simplest resonant configuration, when the Wilton ripple bifurcates from a superposition of $\cos(x)$ and $\cos(2x)$. Among the primary challenges in proving the existence of the remaining Wilton ripples, which bifurcate from a superposition of $\cos(x)$ and $\cos(Kx)$ for any $K\in\mathbb{N}\setminus\{1\}$, is to prove their analyticity with respect to a small-amplitude parameter.
This paper develops a proof that shows for the first time the existence of all Wilton ripple solutions of a nonlinear dispersive equation, namely the Kawahara equation. The proof relies on a carefully constructed Lyapunov-Schmidt reduction to show the analyticity of each solution with respect to a small-amplitude parameter. High-order asymptotic expansions of these solutions, derived in [14], are then used to complete the proof. The author anticipates that this proof can be extended to show the existence of all Wilton ripples of many other nonlinear dispersive equations, including those for gravity-capillary water waves, provided the necessary asymptotic expansions are computed. It is also expected that this proof will connect the classic approach to Wilton ripples, where equation parameters like $\beta$ are fixed, and more contemporary approaches that study sheets of Wilton ripples [8], where the superposition coefficients $a$ and $b$ are instead fixed.

### v2

@@ 2.3.1 Case 1: $K=2$
[…]
@@
$\tilde{\mathcal{B}}(a,b,c_{r})$ $=\begin{pmatrix}ab+ac_{r}-\frac{5}{4}a^{3}-\frac{11}{8}ab^{2}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}\\ \frac{1}{2}a^{2}+bc_{r}-\frac{11}{8}a^{2}b-\frac{91}{72}b^{3}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}\end{pmatrix}=0.$ (54)
@@
[…]
@@
From (56), we have that $\tilde{\tilde{\mathcal{B}}}(0,\pm 1,\pm 1)=0$. The corresponding Jacobian evaluated at $(0,\pm 1,\pm 1)$ is
$\frac{\partial\tilde{\tilde{\mathcal{B}}}}{\partial(\tilde{b},\tilde{c_{r}})}(0,\pm 1,\pm 1)=\begin{pmatrix}\frac{1}{\sqrt{2}}&-\frac{1}{\sqrt{2}}\\ \mp 1&\mp 1\end{pmatrix}$, (57)
which is invertible in both cases. The Analytic Implicit Function Theorem [6, Theorem 2.3] guarantees the existence and uniqueness of
@@
such that, for sufficiently small $a$, we have (i) $\tilde{\tilde{\mathcal{B}}}(a,\tilde{b}^{\pm}(a),\tilde{c_{r}}^{\pm}(a))=0$, (ii) $\tilde{b}^{\pm}(0)=\pm 1$ and $\tilde{c_{r}}^{\pm}(0)=\pm 1$, and (iii) $\tilde{b}^{\pm}(a)$ and $\tilde{c_{r}}^{\pm}(a)$ are real analytic in $a$.
@@
$c(a)=c_{0}-\frac{a}{\sqrt{2}}\tilde{c_{r}}^{\pm}(a)=\frac{4}{5}-\frac{a}{\sqrt{2}}\tilde{c_{r}}^{\pm}(a),$ (59)
@@ 2.3.2 Case 2: $K=3$
$\tilde{\mathcal{B}}(a,b,c_{r})$ $=\begin{pmatrix}ac_{r}-\frac{7}{9}a^{3}+a^{2}b-\frac{34}{63}ab^{2}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}\\ bc_{r}+\frac{1}{3}a^{3}-\frac{34}{63}a^{2}b-\frac{211}{189}b^{3}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}\end{pmatrix}=0.$ (64)
@@
[…]
@@
$\tilde{\tilde{\mathcal{B}}}(a,\tilde{b},\tilde{c_{r}})=\begin{pmatrix}-\frac{7}{9}+\tilde{c_{r}}+\tilde{b}-\frac{34}{63}\tilde{b}^{2}+\mathcal{O}\left(|a|\left(1+|\tilde{b}|+|a|^{2}|\tilde{c_{r}}|\right)^{4}\right)\\ \frac{1}{3}-\frac{34}{63}\tilde{b}+\tilde{b}\tilde{c_{r}}-\frac{211}{189}\tilde{b}^{3}+\mathcal{O}\left(|a|\left(1+|\tilde{b}|+|a|^{2}|\tilde{c_{r}}|\right)^{4}\right)\end{pmatrix}=0.$ (66)
@@
$\tilde{\tilde{\mathcal{B}}}\left(0,\tilde{b}_{0}^{(\sigma)},\tilde{c_{r,0}}^{(\sigma)}\right)=0,$ (70)
@@
$\frac{\partial\tilde{\tilde{\mathcal{B}}}}{\partial(\tilde{b},\tilde{c_{r}})}(0,\tilde{b}_{0}^{(\sigma)},\tilde{c_{r,0}}^{(\sigma)})=\begin{pmatrix}1&1\\ -\frac{34}{63}&\tilde{b}_{0}^{(\sigma)}\end{pmatrix}$, (71)
which are invertible since $\tilde{b}_{0}^{(\sigma)}\neq-34/63$ for all $\sigma$ by contradiction. The Analytic Implicit Function Theorem then gives the existence and uniqueness of
@@
for each $\sigma\in\{1,2,3\}$ and sufficiently small $a$ such that (i) $\tilde{\tilde{\mathcal{B}}}(a,\tilde{b}^{(\sigma)}(a),\tilde{c_{r}}^{(\sigma)}(a))=0$, (ii) $\tilde{b}^{(\sigma)}(0)=\tilde{b}_{0}^{(\sigma)}$ and $\tilde{c_{r}}^{(\sigma)}(0)=\tilde{c_{r,0}}^{(\sigma)}$, and (iii) $\tilde{b}^{(\sigma)}(a)$ and $\tilde{c_{r}}^{(\sigma)}(a)$ are real analytic in $a$.
@@
$c(a)=c_{0}+a^{2}\tilde{c_{r}}^{(\sigma)}(a)=\frac{9}{10}+a^{2}\tilde{c_{r}}^{(\sigma)}(a),$ (73)
@@ 2.3.3 Case 3: $K\geq 4$
$\tilde{\mathcal{B}}(a,b,c_{r})$ $=\begin{pmatrix}ac_{r}+v_{3,0,0}(K)a^{3}+v_{1,2,0}(K)ab^{2}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}\\ bc_{r}+w_{2,1,0}(K)a^{2}b+w_{0,3,0}(K)b^{3}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}\end{pmatrix}=0.$ (79)
@@
$\tilde{\tilde{\mathcal{B}}}(a,\tilde{b},\tilde{c_{r}})=\begin{pmatrix}v_{3,0,0}(K)+\tilde{c_{r}}+v_{1,2,0}(K)\tilde{b}^{2}+\mathcal{O}\left(|a|\left(1+|\tilde{b}|+|a|^{2}|\tilde{c_{r}}|\right)^{4}\right)\\ w_{2,1,0}(K)\tilde{b}+\tilde{c_{r}}\tilde{b}+w_{0,3,0}(K)\tilde{b}^{3}+\mathcal{O}\left(|a|\left(1+|\tilde{b}|+|a|^{2}|\tilde{c_{r}}|\right)^{4}\right)\end{pmatrix}=0.$ (80)
@@
$\tilde{\tilde{\mathcal{B}}}(0,0,-v_{3,0,0}(K))=0,$ (83)
@@
$\frac{\partial\tilde{\tilde{\mathcal{B}}}}{\partial(\tilde{b},\tilde{c_{r}})}(0,0,-v_{3,0,0}(K))=\begin{pmatrix}0&1\\ w_{2,1,0}(K)-v_{3,0,0}(K)&0\end{pmatrix}$, (84)
@@
Thus, (84) is invertible, and the implicit function theorem gives us real analytic functions $\tilde{b}(a)$ and $\tilde{c_{r}}(a)$ such that $\tilde{\tilde{\mathcal{B}}}(a,\tilde{b}(a),\tilde{c_{r}}(a))=0$, $\tilde{b}(0)=0$, and $\tilde{c_{r}}(0)=-v_{3,0,0}(K)$.
@@
$c(a)=c_{0}+a^{2}\tilde{c_{r}}(a)=\frac{K^{2}}{K^{2}+1}+a^{2}\tilde{c_{r}}(a),$ (86)
@@
Unlike before, $\tilde{b}(0)=0$ when $K\geq 4$. As a consequence, it is unclear whether $\tilde{b}(a)$ vanishes identically for sufficiently small $a$. If this is the case, then our Wilton ripple solutions degenerate into Stokes wave solutions with profiles that are orthogonal to $\cos(Kx)$. In the following lemma, we prove that this degeneracy is, in fact, not possible. An important step of our argument appeals to [17], where formal asymptotic computations of (86)-(87) were carried out to arbitrarily high order in $a$ via a Poincaré-Lindstedt method.
Lemma 1.
For each $K\geq 4$, $\tilde{b}(a)\not\equiv 0$ for sufficiently small $a$.
Proof.
Fix $K\geq 4$. Since $\tilde{b}(a)$ and $\tilde{c}(a)$ are real analytic for sufficiently small $a$ and since
$u_{r}(x;a,a\tilde{b}(a),a^{2}\tilde{c}_{r}(a))$ is a composition of real analytic functions in its final three arguments, $c(a)$ and $u(x;a)$ are real analytic for sufficiently small $a$ with values in $\mathbb{R}$ and $\textrm{H}^{4}_{\textrm{per,even}}(\mathbb{T})$, respectively. Hence, for sufficiently small $a$, we have unique power series
$c(a)$ $=\sum_{j=0}^{\infty}c_{j}a^{j},\quad\textrm{absolutely convergent in}~\mathbb{R},$ (89a)
$u(x;a)$ $=\sum_{j=1}^{\infty}u_{j}(x)a^{j},\quad\textrm{absolutely convergent in H}^{4}_{\textrm{per,even}}(\mathbb{T}),$ (89b)
with $c_{j}\in\mathbb{R}$ and $u_{j}(x)\in\textrm{H}^{4}_{\textrm{per,even}}(\mathbb{T})$. By (86),
$c_{0}=\frac{K^{2}}{K^{2}+1}\quad\textrm{and}\quad c_{2}=\frac{(K^{2}+1)(5K^{2}-24)}{6K^{2}(K^{2}-4)}.$ (90)
[…]
$u_{1}(x)=\cos(x).$
Explicit expressions of $c_{j}$ and $u_{j}(x)$ for larger values of $j$ require further expansions of the auxiliary and bifurcation equations.
Returning to the original traveling-wave equation (21), we have
$F(u(x;a),c(a))=0,\quad\textrm{where}$ (91)
$F(u,c):=(c+\mathcal{L})u+u^{2},\quad\mathcal{L}:=\partial_{x}^{2}+\beta\partial_{x}^{4},\quad\beta=1/(1+K^{2}),\quad\textrm{and}\quad K\geq 4.$ (92)
Since $c(a)$ and $u(x;a)$ are real analytic for sufficiently small $a$, $F(u(x;a),c(a))$ is also real analytic with values in $\textrm{L}^{2}_{\textrm{per,even}}(\mathbb{T})$. Thus, $F(u(x;a),c(a))$ admits a unique power series in $a$. Substituting expansions (89a)-(89b) into (91) and using the Cauchy product to rearrange terms as needed [^10] Rearrangement is permissible since $c(a)$ and $u(x;a)$ have absolutely convergent power series and since $F(u,c)$ is a real analytic map of $u$ and $c$ into $\textrm{L}^{2}_{\textrm{per,even}}(\mathbb{T})$. yields the following power series of $F(u(x;a),c(a))$:
$F(u(x;a),c(a))=\sum_{j=1}^{\infty}\biggr(\sum_{i=1}^{j}c_{j-i}u_{i}(x)+\mathcal{L}u_{j}(x)+\sum_{i=1}^{j-1}u_{j-i}(x)u_{i}(x)\biggr)a^{j},$ (93)
which is absolutely convergent in $\textrm{L}^{2}_{\textrm{per,even}}(\mathbb{T})$ for sufficiently small $a$. From (91) and (93), we have
$\big(c_{0}+\mathcal{L}\big)u_{j}(x)=-\sum_{i=1}^{j-1}\Big(c_{j-i}+u_{j-i}(x)\Big)u_{i}(x)\quad\textrm{for}\quad j\in\mathbb{N},$ (94)
where, for notational convenience, the sum on the right-hand side collapses when $j=1$.
In addition to (91), we also have the following for $\omega\in\{1,K\}$:
$\mathcal{F}_{\omega}\Big[F(u(x;a),c(a))\Big]$ $=0,$
$\mathcal{F}_{\omega}\left[c(a)u(x;a)+u(x;a)^{2}\right]+\mathcal{F}_{\omega}\Big[\mathcal{L}u(x;a)\Big]$ $=0,$
$\mathcal{F}_{\omega}\Big[c(a)u(x;a)+u(x;a)^{2}\Big]-c_{0}\mathcal{F}_{\omega}\Big[u(x;a)\Big]$ $=0,$
$\mathcal{F}_{\omega}\Big[\Big(c(a)-c_{0}+u(x;a)\Big)u(x;a)\Big]$ $=0,$ (95)
where the third equality follows from the self-adjointness of $\mathcal{L}$ and the definition of $c_{0}$. The left-hand side of (95) is real analytic for sufficiently small $a$ with values in $\mathbb{R}$. Substituting expansions (89a)-(89b) into (95) yields the unique power series
$\sum_{j=2}^{\infty}\mathcal{F}_{\omega}\Big[\sum_{i=1}^{j-1}\big(c_{j-i}+u_{j-i}(x)\big)u_{i}(x)\Big]a^{j}=0,$ (96)
which converges absolutely in $\mathbb{R}$ for sufficiently small $a$. It follows that
$\mathcal{F}_{\omega}\Big[\sum_{i=1}^{j-1}\big(c_{j-i}+u_{j-i}(x)\big)u_{i}(x)\Big]=0\quad\textrm{for}\quad j\in\mathbb{N}\setminus\{1\}\quad\textrm{and}\quad\omega\in\{1,K\}.$ (97)
Together, (94) and (97) are (up to notation) exactly the hierarchy of equations studied in the formal work of [17] to obtain arbitrarily high-order asymptotic expansions of $c(a)$ and $u(x;a)$ via the Poincaré-Lindstedt method. We have hence demonstrated that these formal computations are consistent with the functional-analytic framework developed in this paper. That is, the formal, high-order asymptotic expansions computed by [17] converge for sufficiently small $a$ to the $c(a)$ and $u(x;a)$ obtained in this work. Having established the rigorous validity of these expansions, we turn to [17, Lemma 5], which shows $\mathcal{F}_{K}[u(x;a)]\not=0$ and, more specifically, $\mathcal{F}_{K}[u(x;a)]=\mathcal{O}\left(a^{K-2}\right)$ as $a\rightarrow 0$. Consequently, $\tilde{b}(a)\not\equiv 0$ for sufficiently small $a$ and, in fact, $\tilde{b}(a)=\mathcal{O}\left(a^{K-3}\right)$ as $a\rightarrow 0$. ∎
With Lemma 1 in hand, we conclude for each $K\geq 4$ that (87) represents a non-degenerate Wilton ripple solution depending real analytically on $a$ with values in $\textrm{H}^{4}_{\textrm{per,even}}(\mathbb{T})$. As in the previous cases, we apply a bootstrap argument to (5) to promote $u(x;a)$ to the more regular space $\textrm{H}^{\infty}_{\textrm{per,even}}(\mathbb{T})$, completing the proof of Theorem 1.
Conclusion
In this paper, we have rigorously established the existence of all 1:K1\!:\!K Wilton ripple solutions of the Kawahara equation for each $K\in\mathbb{N}\setminus\{1\}$. These solutions are defined as small-amplitude, periodic, traveling-wave solutions that bifurcate from a linear combination of $\cos(x)$ and $\cos(Kx)$ in the limit of vanishing amplitude, where the Bond parameter $\beta=1/(1+K^{2})$ induces a degeneracy in the linearized dispersion relation. The resulting bifurcation is of codimension one and leads to a two-dimensional kernel, making the construction of these solutions substantially more delicate than for standard Stokes waves.
Our proof proceeds via a Lyapunov–Schmidt reduction, yielding a reduced bifurcation equation with real analytic dependence on a small-amplitude parameter. The central technical challenge is to verify that the coefficient function $b(a)$ in the Wilton ripple expansion (8) does not vanish identically. This requires a precise understanding of the asymptotic structure of the Wilton ripple solutions and their dependence on $K$. By rigorously justifying the high-order formal expansions developed in [17], we establish the nontriviality of $b(a)$ and thereby conclude the existence of genuine $1:K$ Wilton ripple branches for each $K$.
Previous rigorous existence results for Wilton ripples have largely been restricted to the 1:2 resonance case [2, 3, 22, 23] or relied on bifurcation frameworks that allow equation parameters such as the Bond number $\beta$, vorticity, or stratification to vary [9, 10, 16, 19, 29]. While these approaches yield families of multimodal solutions organized in parameter sheets, they do not directly imply the existence of Wilton ripples in the classical sense, where equation parameters are fixed [1]. The results presented here bridge that gap, confirming the existence of discrete Wilton branches at fixed $\beta$ for all $K$, in a setting that admits full analytic control.
The framework developed in this work is expected to generalize to other nonlinear dispersive models, including the full gravity–capillary water wave equations. A major technical hurdle, however, is that, for each $K\in\mathbb{N}\setminus\{1\}$, the Wilton ripple solutions must be expanded to increasingly high orders in the small-amplitude parameter in order to detect the resonant $\cos(Kx)$ mode. While such computations remain out of reach in the full water wave setting, the formal expansions carried out in [17] for weakly nonlinear models raise the intriguing possibility that analogous expansions could eventually be constructed for more physically realistic systems. Should that become possible, the method of proof outlined here may offer the first rigorous existence theory for all 1:K1\!:\!K Wilton ripples in the full gravity–capillary water wave equations. Future work may also explore whether these fixed-parameter branches of Wilton ripples lie embedded within the bifurcation sheets arising in more recent treatments of multimodal solutions.